The useful life of an artificial heart valve is normally distributed with a mean of 44 months and a standard deviation of 3 months. Determine the probability that the valve will last at least 44 month(s), before wearing out. (Find P( X > 44 )
Answer #1

Answer: 50%

Explanation: It is given that,

Mean life = m = 44 months

Standard eviation = SD = 3 months

Z = (Target value - m) / SD

= (44 - 44) / 3 = 0.00

P(Z<=0) = 0.50

Hence the probability that the valve will last at least 44 months before wearing out is

= 1 - P(Z<=0)

= 1 - 0.50

= 0.50